# Obfuscated Location Disclosure for Remote ID Enabled Drones

### v1

@@ Abstract
In this paper, we propose Obfuscated Location disclOsure for RID-enabled drones (OLO-RID), a solution modifying and extending the RID regulation while preserving drones’ location privacy. Rather than disclosing the actual drone’s location, drones equipped with OLO-RID disclose a differentially private obfuscated location in a mobile scenario. OLO-RID also extends RID messages with encrypted location information, accessible only by authorized entities and valuable to obtain the current drone’s location in safety-critical use cases. We design, implement, and deploy OLO-RID on a Raspberry Pi 3 and release the code of our implementation as open-source. We also perform an extensive performance assessment of the runtime overhead of our solution in terms of processing, communication, memory, and energy consumption. We show that OLO-RID can generate RID messages on a constrained device in less than $0.16$ s while also requiring a minimal energy toll on a relevant device ($0.0236\%$ of energy for a DJI Mini 2). We also evaluate the utility of the proposed approach in the context of three reference use cases involving the drones’ location usage, demonstrating minimal performance degradation when trading off location privacy and utility for next-generation RID-compliant drone ecosystems.
@@ 1 Introduction
The adoption of Unmanned Aerial Vehicles, a.k.a. drones, is increasing today in many application areas, e.g., search-and-rescue, surveillance, goods delivery, and agriculture [1]. Such pervasive diffusion is also supported by reports of leading market companies, forecasting over 9.6 million consumer drone unit shipments globally by 2030, with 865 million devices already deployed in 2023 [2][3].
@@
At the same time, relevant regional aviation authorities need to integrate drones into the local airspaces to enhance monitoring and accountability of drone operations. To this aim, many aviation authorities, e.g., the US-based Federal Aviation Administration (FAA) and the EU-based European Union Aviation Safety Agency (EASA), mandated remote identification of drones through dedicated Remote ID (RID) rules [4][5]. According to such regulations, drones should constantly broadcast on the wireless channel RID messages containing a unique identifier, the drone’s current location (latitude, longitude, altitude) and velocity, the location of the corresponding Control Station (CS), a time mark, and an emergency status indication. RID messages should be emitted no less than once a second in plain text, using either WiFi or BLE—when such communication technologies are unavailable onboard, the drone should be equipped with dedicated external RID modules.
@@
Despite solving drones’ accountability issues, RID regulations introduce confidentiality and privacy issues. Adversaries can track drones using a (network of) passive receiver(s). In turn, such tracking can lead to the leakage of private information about the drone, its pilot, and the drone-based business. For instance, the source location can be traced back to the location of warehouses, the flight destination can lead to customers’ private location disclosure, and the track can reveal IP-protected information about the algorithm used for navigation [6]. A few works investigated privacy issues connected with RID, mainly considering the identity of the drone and collision avoidance strategies [6][7]. However, to preserve full compliance with RID regulations, at the time of this writing, no works considered drone’s location privacy. The mentioned problem shares similarities with privacy issues affecting users interacting with Location-Based Services, where users would like to get suggestions based on their current location while not allowing LBSs providers to track them. However, providing location privacy to RID-enabled drones further complicates the problem, due to the required frequent information broadcast that makes the disclosed locations highly correlated in time and space. Moreover, compared to users’ devices (smartphones and laptops), drones may not be persistently connected to the Internet. Also, drones can be much more constrained in terms of processing, memory, communication, and energy availability, calling for very lightweight approaches. Overall, although we know from the literature that any location obfuscation mechanism is bound to cause privacy loss when used repeatedly [8], a solution specifically tailored to the use cases of RID-enabled drones is currently not available.
@@
Contribution. In our work, we design Obfuscated Location disclOsure for RID-enabled drones (OLO-RID), a framework that allows RID-compliant drones to protect as much as possible their current location while allowing relevant location-based services (using such location) to keep working reliably and efficiently. Our solution integrates and extends the mechanism proposed by Xiao et al. in [9], making it suitable to run at runtime on drones while fulfilling all the requirements imposed by RID. We also extend the RID regulation through encrypted location reports that relevant authorities can use to disclose the drone location if necessary. We implement our proposed solution on two Proof-of-Concepts devices, using a general-purpose laptop and a constrained embedded system Raspberry PI 3 Model B+ (RPI), featuring similar hardware constraints to medium-end drones, and release the code of our implementation as open-source [10]. Through an extensive experimental measurement campaign, we show that OLO-RID can run on relevant platforms by requiring less than $0.16$ s and $0.0236\%$ of energy— considering the energy available on a DJI Mini 2. We also introduce three among the most relevant real-world use cases involving the usage of the drones’ location and extract utility metrics to determine the impact of the disclosure of an obfuscated location on the quality of the services offered. Overall, we evaluate the trade-off between drones’ location privacy and utility, showing the possibility of letting the two requirements coexist meaningfully and effectively.
@@
Note that this work extends and completes our preliminary study in [11] through the following contributions.
@@
  • We replace the basic Geo-Indistinguishability method used in [11] with a solution inspired by the proposal by Xiao et al. in [9], specifically conceived for time-correlated location disclosures, modified on purpose to fit the scenario of RID-enabled drones.
@@
  • We implement the proposed solution on two PoC platforms and release the code as open-source [10].
@@
  • We apply our solution to three relevant real-world use cases in the content of drone-based location-based services. We then show the trade-off existing between location privacy and utility in our proposed solution.
@@ 2.1 Remote ID
Remote ID (RID) is a standard regulation first published in 2021 by the FAA [4], whose goal is to enhance the accountability of UAV operations and use the airspace efficiently while prioritizing public safety. This regulation forces almost all UAVs to emit wireless messages at least once each second, reporting the drone’s unique identity and location, the pilot’s (CS) location, the timestamp and emergency status. The regulation defines two RID modes, i.e., Network RID and Broadcast RID. In Network RID, the UAV delivers unicast RID messages to a third-party service provider over the Internet, possibly relaying such messages through the CS. In Broadcast RID, the UAV broadcasts RID messages using Wi-Fi or Bluetooth so that any compatible receiver nearby can receive such information. Drones not equipped with a compatible transceiver must integrate external RID modules provided by relevant companies. This manuscript focuses explicitly on the Broadcast RID mode, as this scenario presents unique challenges in realistic operational conditions. The RID rule applies to drones weighing more than $250$ g anywhere except FAA-Recognized Identification Areas, explicitly listed in the regulation. RID does not specify any security measures applying to the delivered messages, leaving them free to the implementation. Thus, by default, RID messages are delivered in plaintext. Also, the RID regulation does not provide specific techniques for deploying RID in real-world scenarios. The Drone Remote ID Protocol (DRIP) Working Group (WG) of the IETF takes such gaps into account and is currently working on deployment and security requirements connected to RID [12]. Throughout several documents and Request for Comments, the WG defines the network architecture for RID-compliance deployments, requirements and necessary protocols [13][14]. In this work, we align the network architecture, the names of the entities, and the design requirements to the ones defined by DRIP to create a standard-compliant solution and smooth the possible integration of our solution into standardization activities.
@@ 2.2 Public Key Encryption
Public Key Encryption ( Public Key Encryption (PKE)) techniques allow encrypting a plaintext message $m$ using the public key of the recipient $pk_{R}$ to a ciphertext $c$, such that the recipient, in possession of the corresponding private key $sk_{R}$, is the only party able to decrypt the ciphertext and obtain the plaintext.
@@
In this manuscript, we use the well-known Elliptic Curve Integrated Encryption Scheme (ECIES) as a PKE technique [15].
@@ 3.1 Scenario and assumptions
We consider a UAV compliant with the RID regulation. Thus, it integrates a RID Module that wirelessly broadcasts RID messages from takeoff to shutdown. As discussed in Sec. 2.1, RID messages contain a unique identifier of the UAV $UID$, the UAV’s location data (latitude, longitude, and altitude) $\mathbf{z}$, the UAV’s current velocity $\mathbf{v}$, the CS’s current location (latitude, longitude, and altitude) $\mathbf{x_{CS}}$, a timestamp $t$, and the emergency status indication $em_{t}$. As per the regulation, we consider that the UAV has to broadcast a RID message at least once a second, reporting every time an updated location. We consider drones broadcasting RID messages via WiFi, so as to maximize transmission range while keeping compliance with the RID rule. Consider that WiFi transceivers installed on drones can reach extensive ranges, up to 7 km, guaranteeing sufficient coverage [16]. We consider UAVs carrying a Global Navigation Satellite System (GNSS) Module to obtain the current location $\mathbf{x}$, e.g., a Global Positioning System (GPS) receiver. Moreover, we consider UAVs with enough computational power and energy to run basic cryptography algorithms, such as a PKE scheme, through an Encryption Module. Based on the considered use case (see Sec. 3.3), the drone might operate in scenarios with an intermittent or unavailable Internet connection. Therefore, our solution (see Sec. 5) does not require the existence of a persistent connection of the drone to any Cloud Service Provider (CSP). Although some use cases discussed below might describe the connection of drones to external services, such connection is never required to execute our solution. We also envision the presence of a dedicated process, namely the Obfuscation Module, aimed at generating an obfuscated location according to our solution. Onboard the drone, for some use cases, we also consider the presence of a Communication Module, aimed at establishing mutual communication with other entities. This module can be installed by the owner of the drone, based on the envisioned use case. We do not make any assumption about the operational mode of the UAV, which can be either remotely piloted or autonomous. However, as for the CS location, we consider it is either already stored on the UAV or updated thanks to real-time communication with the CS.
@@
We also consider one or more wireless receivers, which we refer to as observers, in line with the standard DRIP architecture. Observers are entities that can receive RID messages on a compatible RF interface. As defined in RFC 9153 [13], we consider multiple types of observers, namely General Public Observers, and Public Safety Observers. Public Safety Observers are part of a more extensive sensing infrastructure focusing on public safety and security. General Public Observers can be regular users receiving RID messages without relationships with specific sensing infrastructures.
@@
Our system model also includes a Trusted Third Party (TTP), responsible for storing identifiers and cryptographic materials used to secure RID messages. As per the architecture defined by DRIP, we refer to the TTP as the Unmanned Service Supplier (USS). Storage and retrieval of information from the USS can be done through dedicated registries. In line with DRIP, we distinguish between public and private registries [12]. Using the drones’ unique identifier, anyone can use the public registry to retrieve public information. Conversely, only authorized parties can use the private registry to retrieve confidential data. The TTP is always available online, so parties can interact with it if required.
@@
In our work, we want to design a location obfuscation mechanism to protect drones’ actual location when disclosing RID messages, despite the presence of time correlation in the disclosed locations. In Sec. 3.3, we provide a few use cases relevant to our investigation. We report the main notation used in the manuscript in the Appendix (supplementary materials).
@@ 3.2 Adversary Model
The adversary considered in this work is a passive eavesdropper, namely $\epsilon$, who aims to track the UAV based on the disclosed RID messages. As drones deliver RID messages in plaintext, $\epsilon$ can achieve this aim by simply using a receiver compatible with the communication technology (WiFi). Using the retrieved location information, the attacker can track the UAV and deduct private information about the business the drone is used for and the pilot. Such information can be further exploited to take a UAV down, capture it and steal any carried goods or private data stored onboard. All such actions could damage the UAV, the business using the drones, and potentially the safety of the people involved. In line with relevant literature [6] [17], our work considers honest UAVs. We intentionally do not consider scenarios where the pilot purposely turns off RID or tamper with the RID module to carry out unauthorized actions, as this would involve malicious UAVs, which are not in the scope of our paper nor the RID regulation. On the contrary, our work intends to protect the privacy of RID-compliant drones and the business using the drone from tracking and other threats, to finally unleash the potential behind drone usage in various application domains.
@@ 3.3 Use Cases
Use Case 1: Unintentional Invasion of No-Fly Zone. We consider a legitimate UAV operating near an No-Fly Zone (NFZ). An NFZ is an area where UAVs are not allowed to fly, e.g., around Critical Infrastructures such as airports and power distribution grids. We consider that the CI operator maintaining the site deploys one or more observers to monitor the area of the NFZ for drone invasions (among other factors)—based on DRIP, such observers are public safety observers. Based on the deployed observers’ area coverage, the system’s overall reception range could extend outside the boundaries of the NFZ. We define the portion of the coverage area outside of the NFZ as the Warning Area (WA). In principle, UAVs are allowed to fly in the WA. When receiving RID messages, the observers use the disclosed location to check if the drone is invading the NFZ. If this is the case, they take action and use the disclosed location to take the drone down (e.g., by jamming or physically shutting it down–through weapons or nets) [18]. Consider drones disclosing an obfuscated location through RID messages rather than the actual one, and approaching the NFZ. Several situations may occur: (i) the disclosed location lies outside of the NFZ, and also the actual location lies outside of the NFZ–this is a True Negative (TN); (ii) the disclosed location is inside of the NFZ, but the actual location lies outside of the NFZ—this is a False Positive (FP); (iii) the disclosed location lies outside of the NFZ, but the actual location is inside the NFZ—this is a False Negative (FN); and finally, (iv) the disclosed location is inside the NFZ, and also real location lies is inside the NFZ—this is a True Positive (TP). Moreover, when detecting a disclosed location inside the NFZ, the CI operator has to know the actual location to verify the invasion and possibly take action. However, the actual location of the UAV has not been disclosed. In turn, such situations would have a significant impact, as UAVs inadvertently invading the NFZ might cause severe safety and operational issues. Therefore, protecting drones’ location privacy while allowing CI providers to detect invasions and take drones down is challenging and requires the design of a tailored solution. We remark that, for this use case, we consider benign drones unintentionally invading the NFZ. Malicious drones, intentionally invading the NFZ to cause damage, cannot be detected by relying only on RID messages, since malicious drone operators might easily stop transmitting RID messages. Conversely, the aim of this use case is to investigate the trade-off between location privacy and utility for benign drones.
@@
Use Case 2: Discovery of the closest Charging Station. Consider a Service Provider (SP) managing a charging support system for Electric Vehicles consisting of a group of Charging Stations. The SP distributes several charging stations in a given area and allows any EVs, including drones, to recharge by paying a fee. We consider that charging stations can receive RID messages and communicate with a central server online. Considering all the charging stations that received RID messages from a given UAV, being the only entity knowing the exact location of the control stations, the central server can determine the nearest control station to the UAV disclosed location. A control station can communicate with UAVs, e.g., through a secure communication link.
@@
Consider a UAV on a mission, implementing location obfuscation to achieve location privacy. During the mission, the UAV needs to find the closest charging station to recharge its battery. As the UAV discloses an obfuscated location through RID messages, the central server of the charging support system can compute the nearest charging station only based on the disclosed location, not the actual one (this is never disclosed). As a result, the system might not be able to suggest to the UAV the charging station closest to the UAV’s actual location. A sub-optimal choice can cause extra distance to be travelled by the drone and more energy consumption, possibly leading to battery drain, energy exhaustion and mission failure. Therefore, protecting drones’ location privacy while allowing SPs to provide service is challenging and requires the design of a tailored solution. We remark that the aim of this use case is to investigate the trade-off between location privacy and utility for drone-based LBS. Real charging stations might also use a different protocol, e.g., broadcasting their own location so to allow UAV to determine the best charging station autonomously. However, since no trade-off would be involved in such a scenario, we do not consider it for our analysis.
@@
Use Case 3: Drone-as-a-Service. We consider drones deployed by a telecommunication SP and providing a service to users nearby, in line with the emerging Drone-as-a-Service (DaaS) scenario [19]. In particular, we assume such drones are equipped with a Subscriber Identity Module (SIM) card that allows them to connect to the mobile cellular network, relaying requests coming from active users. Thus, users without Internet connectivity in remote regions can connect to such UAVs, establish a secure connection, and make phone calls or access services available online.
@@
Consider that the drones implement location obfuscation to achieve location privacy. Thus, when users would like to use a service provided by the SP via the drones, they connect to the Radio Frequency (RF) interface (WiFi), detect all UAVs in a given area and pick the one whose disclosed distance (via RID) is the least from the user location, assuming the closest drone is the one providing the best signal quality. As the UAVs disclose obfuscated locations, the user might make a sub-optimal choice, i.e., pick the UAV whose actual location is not the least. Thus, the signals emitted by the user and the UAV have to travel extra distance than the minimum, with consequent degradation of the achievable Quality of Service (QoS). Therefore, protecting a drone’s location privacy while allowing a SP using such drones to provide reliable and efficient location-based services is challenging and requires the design of a tailored solution.
@@
Note that users of DaaS applications might also use other parameters to pick the best UAV (e.g., the Received Signal Strength-RSS). However, channel quality indicators such as the RSS notoriously exhibit fast-fading phenomena that make distance estimation highly error-prone [20].

### v2

@@ Abstract
 [^1] This is a personal copy of the authors. Not for redistribution. The final version of the paper will be available soon through the digital library IEEExplore at https://ieeexplore.ieee.org/document/11230512/.
@@
In this paper, we propose Obfuscated Location disclOsure for RID-enabled drones (OLO-RID), a solution modifying and extending the RID regulation while preserving drones’ location privacy. Rather than disclosing the actual drone’s location, drones equipped with OLO-RID disclose a differentially private obfuscated location. OLO-RID also extends RID messages with encrypted location information, accessible only by authorized entities and valuable to obtain the current drone’s location in safety-critical use cases. We design, implement, and deploy OLO-RID on a Raspberry Pi 3 and release the code of our implementation as open-source. We also perform an extensive performance assessment of the runtime overhead of our solution in terms of processing, communication, memory, and energy consumption. We show that OLO-RID can generate RID messages on a constrained device in less than $0.16$ s while also requiring a minimal energy toll on a relevant device ($0.0236\%$ of energy for a DJI Mini 2). We also evaluate the utility of the proposed approach in the context of three reference use cases involving the drones’ location usage, demonstrating minimal performance degradation when trading off location privacy and utility for next-generation RID-compliant drone ecosystems.
@@ 1 Introduction
The adoption of Unmanned Aerial Vehicles, a.k.a. drones [^2] In this paper, the terms Unmanned Aerial Vehicles (UAVs) and drones are used interchangeably., is increasing today in many application areas, e.g., search-and-rescue, surveillance, goods delivery, and agriculture [Shakhatreh2019_access]. Such pervasive diffusion is also supported by reports of leading market companies, forecasting over 9.6 million consumer drone unit shipments globally by 2030, with 865 million devices already deployed in 2023 [statista_dronesMarket][statista_dronesUS].
@@
At the same time, relevant regional aviation authorities need to integrate drones into the local airspaces to enhance monitoring and accountability of drone operations. To this aim, many aviation authorities, e.g., the US-based Federal Aviation Administration (FAA) and the EU-based European Union Aviation Safety Agency (EASA), mandated remote identification of drones through dedicated Remote ID (RID) rules [remote_id][belwafi2022_access]. According to such regulations, drones should constantly broadcast on the wireless channel RID messages containing a unique identifier, the drone’s current location (latitude, longitude, altitude) and velocity, the location of the corresponding Control Station (CS), a time mark, and an emergency status indication. RID messages should be emitted no less than once a second in plain text, using either WiFi or BLE—when such communication technologies are unavailable onboard, the drone should be equipped with dedicated external RID modules.
@@
Despite solving drones’ accountability issues, RID regulations introduce confidentiality and privacy issues. Adversaries can track drones using a (network of) passive receiver(s). In turn, such tracking can lead to the leakage of private information about the drone, its pilot, and the drone-based business. For instance, the source location can be traced back to the location of warehouses, the flight destination can lead to customers’ private location disclosure, and the track can reveal IP-protected information about the algorithm used for navigation [wisse2023_iotj]. A real-life example in this direction is the leakage occurred in October 2022, where a massive data leak exposed over 80,000 drone IDs and 90 million drone-monitoring logs from DJI AeroScope devices [aeroscope]. A few works investigated privacy issues connected with RID, mainly considering the identity of the drone and collision avoidance strategies [wisse2023_iotj][tedeschi2022_tdsc]. However, to preserve full compliance with RID regulations, at the time of this writing, no works considered drone’s location privacy. The mentioned problem shares similarities with privacy issues affecting users interacting with Location-Based Services, where users would like to get suggestions based on their current location while not allowing LBSs providers to track them. However, providing location privacy to RID-enabled drones further complicates the problem, due to the required frequent information broadcast that makes the disclosed locations highly correlated in time and space. Moreover, compared to users’ devices (smartphones, laptops), drones may not be persistently connected to the Internet. Also, drones can be more constrained in terms of processing, memory, communication, and energy availability, calling for very lightweight approaches. Overall, although we know from the literature that any location obfuscation mechanism is bound to cause privacy loss when used repeatedly [chatzikokolakis2014_pets], a solution specifically tailored to the use cases of remotely-piloted RID-enabled drones is currently unavailable.
@@
Contribution. In our work, we design Obfuscated Location disclOsure for RID-enabled drones (OLO-RID), a framework that allows RID-compliant remotely-piloted drones to protect as much as possible their current location while allowing relevant location-based services (using such location) to keep working reliably and efficiently. Our solution integrates and extends the mechanism proposed by Xiao et al. in [xiao2015_ccs], making it suitable for running at runtime on drones while fulfilling all the requirements imposed by RID. We also extend the RID regulation through encrypted location reports that relevant authorities can use to disclose the drone location, if necessary. We implement our proposed solution on two Proof-of-Concepts devices, using a general-purpose laptop and a constrained embedded system Raspberry PI 3 Model B+ (RPI), featuring similar hardware constraints to medium-end drones, and release the code of our implementation as open-source [code]. Through an extensive experimental measurement campaign, we show that OLO-RID can run on relevant platforms by requiring less than $0.16$ s and $0.0236\%$ of energy— considering the energy available on a DJI Mini 2. We also introduce three among the most relevant real-world use cases involving the usage of the drones’ location and extract utility metrics to determine the impact of the disclosure of an obfuscated location on the quality of the services offered. Overall, we evaluate the trade-off between drones’ location privacy and utility, showing the possibility of letting the two requirements coexist meaningfully and effectively.
@@
Note that this work extends and completes our preliminary study in [brighente2022_ares] through the following contributions.
@@
  • We replace the basic Geo-Indistinguishability method used in [brighente2022_ares] with a solution inspired by the proposal by Xiao et al. in [xiao2015_ccs], specifically conceived for time-correlated location disclosures, modified on purpose to fit the scenario of RID-enabled drones.
@@
  • We implement the proposed solution on two PoC platforms and release the code as open-source [code].
@@
  • We apply OLO-RID to three relevant real-world use cases in the content of drone-based location-based services. We then show the trade-off existing between location privacy and utility in our proposed solution.
@@ 2.1 Remote ID
Remote ID (RID) is a standard regulation first published in 2021 by the FAA [remote_id], whose goal is to enhance the accountability of UAV operations and use the airspace efficiently while prioritizing public safety. This regulation forces almost all UAVs to emit wireless messages at least once each second, reporting their unique identity and location, the pilot’s (CS) location, the timestamp and emergency status. The regulation defines two RID modes, i.e., Network RID and Broadcast RID. In Network RID, the UAV delivers unicast RID messages to a third-party service provider over the Internet, possibly relaying such messages through the CS. In Broadcast RID, the UAV broadcasts RID messages using Wi-Fi or Bluetooth so that any compatible receiver nearby can receive such information. UAVs not equipped with a compatible transceiver must integrate external RID modules provided by relevant companies. This manuscript focuses explicitly on the Broadcast RID mode, as this scenario presents unique challenges in realistic operational conditions. The RID rule applies to UAVs weighing more than $250$ g anywhere except FAA-Recognized Identification Areas, explicitly listed in the regulation. RID does not specify any security measures applying to the delivered messages, leaving them free to the implementation. Thus, by default, RID messages are delivered in plaintext. Also, RID does not provide specific techniques for deploying RID in real-world scenarios. The Drone Remote ID Protocol (DRIP) Working Group (WG) of the IETF takes such gaps into account and is currently working on deployment and security requirements connected to RID [drip]. Throughout several documents and Request for Comments, the WG defines the network architecture for RID-compliance deployments, requirements and necessary protocols [rfc9153][rfc9434]. In this work, we align the network architecture, entities, and design requirements to the ones defined by DRIP, so creating a standard-compliant solution and smooth the possible integration of our solution into standardization activities.
@@ 2.2 Public Key Encryption
Public Key Encryption (PKE) techniques allow encrypting a plaintext message $m$ using the public key of the recipient $pk_{R}$ to a ciphertext $c$, such that the recipient, in possession of the corresponding private key $sk_{R}$, is the only party able to decrypt the ciphertext and obtain the plaintext.
@@
In this manuscript, we use the well-known Elliptic Curve Integrated Encryption Scheme (ECIES) as a PKE technique [ecies_standard]. Compared to symmetric encryption techniques, public-key encryption schemes like ECIES allow reducing storage requirements and minimizing the secrets involved, thus easing system management.
@@ 3.1 Scenario and assumptions
We consider a Remotely-Piloted Aircraft System (RPAS), compliant with the RID regulation. Thus, it integrates a RID Module that wirelessly broadcasts RID messages from takeoff to shutdown. As discussed in Sec. 2.1, RID messages contain a unique identifier of the UAV $UID$, the UAV’s disclosed location data (latitude, longitude, and altitude) $\mathbf{z}$, the UAV’s current velocity $\mathbf{v}$, the CS’s current location (latitude, longitude, and altitude) $\mathbf{x_{CS}}$, a timestamp $t$, and the emergency status indication $em_{t}$. As per the regulation, we consider that the UAV has to broadcast a RID message at least once a second, reporting every time an updated location. We consider UAVs broadcasting RID messages via WiFi, so as to maximize transmission range while keeping compliance with the RID rule. Consider that WiFi transceivers installed on UAVs can reach a range of a few kilometers, guaranteeing sufficient coverage [sciancalepore2022_acsac]. We consider UAVs carrying a Global Navigation Satellite System (GNSS) Module to obtain the current location $\mathbf{x}$, e.g., a Global Positioning System (GPS) receiver. Moreover, we consider UAVs with enough computational power and energy to run basic cryptography algorithms, such as a PKE scheme, through an Encryption Module. Based on the considered use case (see Sec. 3.3), the UAV might operate in scenarios with an intermittent or unavailable Internet connection. Therefore, our solution (see Sec. 5) does not require the existence of a persistent connection of the UAV to any Cloud Service Provider (CSP). Although some use cases discussed below might describe the connection of UAVs to external services, such connection is never required to execute our solution. We also envision the presence of a dedicated process, namely the Obfuscation Module, aimed at generating an obfuscated location according to our solution. Onboard the UAV, for some use cases, we also consider the presence of a Communication Module, aimed at establishing mutual communication with other entities. This module can be installed by the owner of the UAV, based on the envisioned use case. As for the CS location, we consider it is either already stored on the UAV or updated thanks to real-time communication with the CS.
@@
We also consider one or more wireless receivers, which we refer to as observers, in line with the standard DRIP architecture. Observers are entities that can receive RID messages on a compatible RF interface. As defined in RFC 9153 [rfc9153], we consider multiple types of observers, namely General Public Observers, and Public Safety Observers. Public Safety Observers are part of a more extensive sensing infrastructure focusing on public safety and security. General Public Observers can be regular users receiving RID messages without relationships with specific sensing infrastructures.
@@
Our system model also includes a Trusted Third Party (TTP), responsible for storing identifiers and cryptographic materials used to secure RID messages. As per the architecture defined by DRIP, we refer to the TTP as the Unmanned Service Supplier (USS). Storage and retrieval of information from the USS can be done through dedicated registries. In line with DRIP, we distinguish between public and private registries [drip]. Using the UAVs’ unique identifier, anyone can use the public registry to retrieve public information. Conversely, only authorized parties can use the private registry to retrieve confidential data. The TTP is always available online, so parties can interact with it if required.
@@
In our work, we want to design a location obfuscation mechanism to protect UAVs’ actual location when disclosing RID messages, despite the presence of time correlation in the disclosed locations. In Sec. 3.3, we provide a few use cases relevant to our investigation. We report the main notation used in the manuscript in the Appendix (supplementary materials).
@@ 3.2 Adversary Model
The adversary considered in this work is a static passive eavesdropper, namely $\epsilon$, who aims to track the UAV based on the disclosed RID messages. As UAVs deliver RID messages in plaintext, $\epsilon$ can achieve this aim by simply using a receiver compatible with the communication technology (WiFi). Using the retrieved location information, the attacker can track the UAV and deduct private information about the business the UAV is used for and the pilot. Such information can be further exploited to take a UAV down, capture it and steal any carried goods or private data stored onboard. All such actions could damage the UAV, the business using the UAVs, and potentially the safety of the people involved. In line with relevant literature [wisse2023_iotj] [tedeschi2023_acsac], our work considers honest UAVs. We intentionally do not consider scenarios where the pilot purposely turns off RID or tamper with the RID module to carry out unauthorized actions, as this would involve malicious UAVs, which are not in the scope of our paper nor in that of the RID regulation. On the contrary, our work intends to protect the privacy of RID-compliant UAVs and the business using the UAV from tracking and other threats, to finally unleash the potential behind their usage in various application domains.
@@ 3.3 Use Cases
Use Case 1: Unintentional Invasion of No-Fly Zone. We consider a legitimate UAV operating near an No-Fly Zone (NFZ). An NFZ is an area where UAVs are not allowed to fly, e.g., around Critical Infrastructures such as airports and power distribution grids. We consider that the CI operator maintaining the site deploys one or more observers to monitor the area of the NFZ for UAV invasions (among other factors)—based on DRIP, such observers are public safety observers. Based on the deployed observers’ area coverage, the system’s overall reception range could extend outside the boundaries of the NFZ. We define the portion of the coverage area outside of the NFZ as the Warning Area (WA). In principle, UAVs are allowed to fly in the WA. When receiving RID messages, the observers use the disclosed location to check if the UAV is invading the NFZ. If this is the case, they take action and use the disclosed location to take the UAV down (e.g., by jamming or physically shutting it down–through weapons or nets) [sciancalepore2023_iotj]. Consider UAVs disclosing an obfuscated location through RID messages rather than the actual one, and approaching the NFZ. Several situations may occur: (i) the disclosed location lies outside of the NFZ, and also the actual location lies outside of the NFZ–this is a True Negative (TN); (ii) the disclosed location is inside of the NFZ, but the actual location lies outside of the NFZ—this is a False Positive (FP); (iii) the disclosed location lies outside of the NFZ, but the actual location is inside the NFZ—this is a False Negative (FN); and finally, (iv) the disclosed location is inside the NFZ, and also real location lies is inside the NFZ—this is a True Positive (TP). Moreover, when detecting a disclosed location inside the NFZ, the CI operator has to know the actual location to verify the invasion and possibly take action. However, the actual location of the UAV has not been disclosed. In turn, such situations would have a significant impact, as UAVs inadvertently invading the NFZ might cause severe security, safety and operational issues [wigchert2025_comnet]. Therefore, protecting UAVs’ location privacy while allowing CI providers to detect invasions and take UAVs down is challenging and requires the design of a tailored solution. We remark that, for this use case, we consider benign UAVs unintentionally invading the NFZ. Malicious UAVs, intentionally invading the NFZ to cause damage, cannot be detected by relying only on RID messages, since malicious UAV operators might easily stop transmitting RID messages. Conversely, the aim of this use case is to investigate the trade-off between location privacy and utility for benign UAVs.
@@
Use Case 2: Discovery of the closest Charging Station. Consider a Service Provider (SP) managing a charging support system for Electric Vehicles consisting of a group of Charging Stations. The SP distributes several charging stations in a given area and allows any EVs, including UAVs, to recharge by paying a fee. We consider that charging stations can receive RID messages and communicate with a central server online. Considering all the charging stations that received RID messages from a given UAV, being the only entity knowing the exact location of the control stations, the central server can determine the nearest control station to the UAV disclosed location. A control station can communicate with UAVs, e.g., through a secure communication link.
@@
Consider a UAV on a mission, implementing location obfuscation to achieve location privacy. During the mission, the UAV needs to find the closest charging station to recharge its battery. As the UAV discloses an obfuscated location through RID messages, the central server of the charging support system can compute the nearest charging station only based on the disclosed location, not the actual one (this is never disclosed). As a result, the system might not be able to suggest to the UAV the charging station closest to the UAV’s actual location. A sub-optimal choice can cause extra distance to be travelled by the UAV and more energy consumption, possibly leading to battery drain, energy exhaustion and mission failure. Therefore, protecting UAVs’ location privacy while allowing SPs to provide service is challenging and requires the design of a tailored solution. We remark that the aim of this use case is to investigate the trade-off between location privacy and utility for UAV-based LBS. Real charging stations might also use a different protocol, e.g., broadcasting their own location so to allow UAV to determine the best charging station autonomously. However, since no trade-off would be involved in such a scenario, we do not consider it for our analysis.
@@
Use Case 3: UAV-as-a-Service. We consider UAVs deployed by a telecommunication SP and providing a service to users nearby, in line with the emerging Drone-as-a-Service (DaaS) scenario [george2023_sacmat]. In particular, we assume such UAVs are equipped with a Subscriber Identity Module (SIM) card that allows them to connect to the mobile cellular network, relaying requests coming from active users. Thus, users without Internet connectivity in remote regions can connect to such UAVs, establish a secure connection, and make phone calls or access services available online.
@@
Consider that the UAVs implement location obfuscation to achieve location privacy. Thus, when users would like to use a service provided by the SP via the UAVs, they connect to the Radio Frequency (RF) interface (WiFi), detect all UAVs in a given area and pick the one whose disclosed distance (via RID) is the least from the user location, assuming the closest UAV is the one providing the best signal quality. As the UAVs disclose obfuscated locations, the user might make a sub-optimal choice, i.e., pick the UAV whose actual location is not the least. Thus, the signals emitted by the user and the UAV have to travel extra distance than the minimum, with consequent degradation of the achievable Quality of Service (QoS). Therefore, protecting a UAV’s location privacy while allowing a SP using such UAVs to provide reliable and efficient location-based services is challenging and requires the design of a tailored solution.
@@
Note that users of DaaS applications might also use other parameters to pick the best UAV (e.g., the Received Signal Strength-RSS). However, channel quality indicators such as the RSS notoriously exhibit fast-fading phenomena that make distance estimation highly error-prone [tedeschi2021_spaccs].